Find the slope of a line perpendicular to the line that passes through the following points (3,9) and (7,15)
Answer by lwsshak3(11628)

Find the slope of a line perpendicular to the line that passes through the following points (3,9) and (7,15)
**
slope of line that passes through points (3,9) and (7,15)=∆y/∆x=rise/run=change of y/change of x
=(15-9)/(7-3)=6/4=3/2
..
slope of line perpendicular to line above=its negative reciprocal=-2/3